The energy from the combustion of Hydrazine, N2H4, is used to power rockets into space in the reaction:


N2H4(g)+O2(g)=N2(g)+2H2O(l) delta H degree= -627.6KJ


How many killograms of hydrazine would be necessary to produce 1.0 X 10^8 KJ of energy?


Hint: One mole of N2H4 produces 627.6KJ of energy. How many moles (and then grams) are required to produce 1.0 X 10^8 KJ of energy.


1
Expert's answer
2021-06-13T11:30:01-0400

"n=\frac{1.0 \times 10^8}{627.6} = 1.5933 \times 10^5 \;mol"

MM(N2H4)= 32.04 g/mol

"m = 1.5933 \times 10^5 \times 32.04 = 51.05 \times 10^5 \;g

= 51.05 \times 10^2 \;kg \\

= 5105 \;kg"

Answer: 5105 kg